# Fundamentals and Advances in Transverse Thermoelectrics

### v1

@@ V.2 Effect of thermal boundary conditions
Figure 7: (a),(b) Schematics of the measurement setup for the transverse thermoelectric generation in the isothermal and adiabatic conditions. (c) Schematic of the measurement setup for ANE in SmCo5-type permanent magnet slabs with different widths $d$. (d) $H$ dependence of $V_{y}/d$ for the SmCo5 slabs with a width of $d=7.7~{}{\rm mm}$ (red) and 2.1 mm (black) at $|\nabla_{x}T|\sim 1\times 10^{3}~{}{\rm K/m}$. (e) Dependence of the ANE-induced electric field $V_{\rm ANE}/d$ on $|\nabla_{x}T|$.
@@
We now experimentally demonstrate the influence of sample dimensions on the ANE-induced transverse thermopower using a bulk polycrystalline SmCo5 slab. The SmCo5 block, commercially available from Magfine Corporation, Japan, was cut into two rectangular slabs with dimensions along the $x$, $y$, and $z$ directions respectively being 10.0, 7.7, and 0.8 mm and 10.0, 2.1, and 0.8 mm using a diamond wire saw, where only the length along the $y$ direction, $d$, is significantly different. The slabs have an easy axis of magnetization along $z$ direction. The measurement configuration is illustrated in Fig. 7(c). The two SmCo5 slabs were bridged between two heat baths separated by 8.0 mm and firmly attached to the heat baths. To apply a temperature difference between the heat baths, chip heaters were attached to the hot side. The sample holder was fixed at the center of a superconducting magnet, where a magnetic field ranging from $\mu_{0}H=5~{}{\rm T}$ to $-5~{}{\rm T}$ was applied along the $z$ direction. The $H$ dependence of the transverse voltage $V_{y}$ along the $y$ direction was recorded for both the slabs at various values of the temperature differences. To accurately measure the temperature gradient on the SmCo5 slabs, the top surface of the slabs was coated with insulating black ink with an emissivity of $>0.94$ and steady-state temperature images were captured using an infrared camera.
@@
Figure 7(d) shows the $H$ dependence of the transverse electric field $V_{y}/d$ for the SmCo5 slabs, recorded under a longitudinal temperature gradient of $|\nabla_{x}T|\sim 1\times 10^{3}~{}{\rm K/m}$. In both the slabs, the transverse electric field changed its sign with large hysteresis with respect to the $H$ reversal. The $H$ dependence of $V_{y}/d$ corresponds to the magnetization curve of the SmCo5 permanent magnet, indicating the appearance of ANE. To determine the magnitude of the ANE-induced transverse thermopower for the SmCo5 slabs, the $H$-odd-dependent component of $V_{y}/d$ under the saturation state, $V_{\rm ANE}/d$, was extracted for each $|\nabla_{x}T|$ value [Fig. 7(e)]. Although the SmCo5 slabs were cut from the same ingot, they exhibited different values of the transverse thermopower. The $V_{\rm ANE}/d$ value for the SmCo5 slab with $d=2.1~{}{\rm mm}$ ($d=7.7~{}{\rm mm}$) was found to be $4.8~{}\mu{\rm V/K}$ ($4.0~{}\mu{\rm V/K}$), where the difference is much larger than measurement errors. Recall the fact that the slab with $d=2.1~{}{\rm mm}$ ($d=7.7~{}{\rm mm}$) is closer to the adiabatic (isothermal) condition. Therefore, the transverse thermopower under the adiabatic condition is expected to be overestimated compared to that under the isothermal condition because of the same sign of $S_{xy}$ and the correction term. This expectation is consistent with the observed difference in the transverse thermopower. The experimental results shown here offer guidelines for determining the appropriate sample aspect ratio and measurement setup for transverse thermopower measurements, minimizing parasitic signals from the Seebeck effect and THE. We also note that considering typical sample dimensions, the estimation of $S_{\rm ANE}$ from the measurement of the anomalous Ettingshausen effect and the Onsager reciprocal relation reduces the influence of the correction term [29].
@@ VI.1 Heat flux sensing
A heat flux sensor is a device enabling simultaneous detection of the magnitude and direction of a heat flow, which can be an essential component for smart thermal management systems. However, practical applications of heat flux sensors have been limited due to several problems. The commercially available sensors are based on the Seebeck effect and consist of a serially connected three-dimensional array of two different thermoelectric materials. Therefore, the heat flux sensors based on the Seebeck effect require a durable substrate or thick plate to provide mechanical stability. Hence, the conventional sensors are mainly applicable to flat surfaces and their mechanical flexibility is limited. Due to the presence of thick substrates, these sensors have large thermal resistance that changes the heat flow distribution to be detected. Although the sensibility of the heat flux sensors based on the Seebeck effect is proportional to the number of the thermoelectric material junctions and to the sensor size, the complex structure and low mechanical durability make it difficult to construct highly sensitive sensors. As an alternative to the sensors having these drawbacks, thin-film-based heat flux sensors driven by the transverse thermoelectric effects, e.g. spin-Seebeck-driven ISHE [149] and ANE [34, 45, 150, 151], were proposed and demonstrated. Owing to its simpler structure and improved transverse thermopower, ANE-based heat flux sensors have recently become mainstream. In the ANE-based sensor, a lateral thermopile structure consisting of alternately arranged and serially connected two different wires with different $S_{\rm ANE}$ values is used. The simple thermopile structure and the symmetry of ANE make the sensors thin, reducing their thermal resistance. If magnetic materials showing large ANE can be formed on thin flexible sheets, flexible heat flux sensors can be constructed [45, 150]. These features of the ANE-based heat flux sensors are advantageous over the conventional heat flux sensors based on the Seebeck effect.
Figure 8: (a) Schematic of the ANE-based heat flux sensor. The sensor comprises the amorphous Sm20Co80 wires and Au wires arranged alternately and connected in series formed on a flexible polyethylene naphthalate (PEN) sheet. ${\bf q}$ denotes the heat flux density with the magnitude $q$. (b) $H$ dependence of $V$ for the Sm20Co80/Au sensor for various values of $q$. (c) $q$ dependence of $V_{\rm ANE}$ at $|\mu_{0}H|=0.2$ and 0.0 T. The data in (b) and (c) are reconstructed from Fig. 6 of Ref. [150].
@@
Here, we show a proof-of-concept demonstration of a flexible heat flux sensor based on ANE. We deposited the 100-nm-thick amorphous Sm20Co80 film on a 50-$\mu$m-thick flexible polyethylene naphthalate (PEN) sheet, where amorphous Sm20Co80 exhibits substantial $S_{\rm ANE}$, coercivity, and remanent magnetization and can be formed on any substrate [150]. We formed 50 Sm20Co80 wires with a width of 50 $\mu$m and a length of 10 mm arranged in a parallel configuration at an interval of 150 $\mu$m. The wires were then connected in series using Au wires to make the total device area of $10\times 10~{}{\rm mm}^{2}$, where $q$ was estimated through the normalization by this area. The total length of the Sm20Co80 wires for this configuration reaches 500 mm. We measured the voltage $V$ between the ends of the thermopile structure with applying an in-plane external magnetic field along the width direction of the wires and an out-of-plane heat current. The details of the experimental procedures are shown in Ref. [150].
As shown in Fig. 8(b), clear ANE signals with hysteresis behaviors were observed in the $H$ dependence of $V$ and their magnitude increases with increasing $q$ across the device. Figure 8(c) shows the ANE voltage $V_{\rm ANE}$ as a function of $q$ [150]. Here, $V_{\rm ANE}$ at 0.2 T was extracted by the field-odd component of $V$ at $|\mu_{0}H|=0.2~{}{\rm T}$ and $V_{\rm ANE}$ at zero field was extracted from the zero-field $V$ values for the positive-to-negative and negative-to-positive sweep of $H$. The sensitivity of the ANE-based heat flux sensor was estimated to be $V_{\rm ANE}/j_{\rm q}=1.12\pm 0.02\times 10^{-7}~{}{\rm VW}^{-1}{\rm m}^{2}$ and $0.89\pm 0.02\times 10^{-7}~{}{\rm VW}^{-1}{\rm m}^{2}$ in the saturation magnetization (0.2 T) and remanent magnetization (0.0 T) states, respectively. Importantly, the use of magnetic thin films with finite coercivity and remanent magnetization makes it possible to sense a heat flux using ANE even in the absence of an external magnetic field. Therefore, the design and control of magnetic anisotropy are essential for the application of ANE. Although the sensitivity of the sensor used in this demonstration is lower than that of commercial sensors based on the Seebeck effect, ANE has strong advantages in flexibility and low thermal resistance, extending applications of heat flux sensors. To further improve the performance of the ANE-based heat flux sensors, extensive efforts have been devoted to developing new materials with larger $S_{\rm ANE}$ and lower $\kappa$ as well as to optimizing the device structure. Since amorphous materials can be fabricated on any surface and have lower $\kappa$ than crystalline materials, they are promising candidates for the development of low-cost, high-performance heat flux sensors with large areas [56, 57, 58, 59, 150].
@@ VI.2 Thermal energy harvesting
Transverse thermoelectric devices are useful for harvesting and utilizing widely distributed thermal energy. By installing simple block- or sheet-shaped transverse thermoelectric devices on a heat source surface, one can extract a charge current and voltage along the surface direction from a temperature gradient perpendicular to the surface. Kirihara et al. pointed out that spin-Seebeck-driven ISHE is suitable for the thermal energy harvesting from curved or complex-shaped heat source surfaces and demonstrated that SSE devices can be fabricated using simple coating [82] or spray [149] processes. Similar functionality can be obtained using ONE (ANE) in the in-plane magnetic field (magnetized) configuration, enabling the fabrication of flexible transverse thermoelectric devices based on thin films deposited on flexible substrates [150] or amorphous metal ribbons [56, 57, 58, 59]. However, challenges include the difficulty of extracting a large charge current and electrical power from thin films and the fact that the thermopower generated by ONE, ANE, and spin-Seebeck-driven ISHE has not yet reached sufficient levels for practical applications.
@@
Experimental demonstrations of thermal energy harvesting using the off-diagonal Seebeck effect, which exhibits high transverse thermoelectric conversion performance, have also been reported. Figure 9 compares the maximum power density normalized by the square of the temperature gradient ($\propto$ the power factor) for thermoelectric conversion modules driven by various principles [4, 9, 10, 158, 159, 160, 161, 91, 92, 93, 156, 100]. The maximum power density of the ATML-based modules driven by the off-diagonal Seebeck effect stands out, being 2-3 orders of magnitude larger than that of the modules driven by ONE and ANE. Recently developed SmCo5/BST ATML exhibits excellent transverse thermoelectric conversion performance, with the maximum power density of the modules composed of 14 ATML elements reaching 56.7 mW/cm2 at a temperature difference of 152 K [100]. SmCo5/BST ATML has higher thermal conductivity than longitudinal thermoelectric materials, making it difficult to maintain a large temperature difference in steady-state conditions; the efficiency of the ATML-based module falls short of that of longitudinal modules driven by the Seebeck effect. However, the maximum power density normalized by the square of the temperature gradient in ATMLs exceeds that of the longitudinal modules based on the Seebeck effect (Fig. 9). Therefore, ATML-based transverse thermoelectric modules are suitable for thermal energy harvesting because they can generate large power even in environments where only small temperature differences can be generated. Additionally, in transverse thermoelectric modules with thermopile structures, by alternately changing the height of each element, a built-in heat sink can be incorporated to promote heat release to the atmosphere [100]. Such a built-in heat sink structure cannot be achieved in longitudinal modules where rigid plates are present on the top and bottom surfaces. Even without sandwiching the thermoelectric module between hot and cold heat baths, the built-in heat sink structure enables substantially large power generation simply by placing it on a heat source owing to the improved heat release to the atmosphere. In particular, SmCo5/BST ATML has remanent magnetization and functions as a permanent magnet; when it can be attached to a heat source through a magnetic attractive force, the contact thermal resistance between the module and heat source decreases, enabling the effective generation of a temperature difference from the heat source. As shown in Ref. [100], when the thermopile module comprising SmCo5/BST ATML elements is placed on a hot plate, the electrical power generated solely by heat release to the air reaches a value sufficient to drive small sensors.
@@ VI.3 Hybrid transverse magneto-thermoelectric conversion
In this subsection, we introduce one of the promising strategies for improving transverse thermoelectric conversion performance: hybrid transverse magneto-thermoelectric conversion [98, 99]. As classified in Fig. 3, there are various driving principles for the transverse thermoelectric conversion, but the phenomena with time-reversal symmetry breaking and those with structural symmetry breaking have been studied independently in different fields. The interdisciplinary fusion of these phenomena not only leads to the creation of new research fields but also significantly contributes to improving the transverse thermoelectric conversion performance. Here, imagine a situation where multiple transverse thermoelectric conversion principles manifest in a single material, i.e. either a homogeneous or composite/hybrid material. As an example, if the transverse thermopower due to the off-diagonal Seebeck effect $S_{\rm OD}$ and that due to ONE $S_{\rm ONE}$ are simultaneously manifested, the total transverse thermopower is simply the sum of these two contribution: $S_{\rm OD}+S_{\rm ONE}$. Since $|S_{\rm OD}|>|S_{\rm ONE}|$ in many cases, the contribution of ONE to the total transverse thermopower is small. In this situation, the figure of merit is proportional to the square of the total transverse thermopower, and described as
@@
The effectiveness of the hybrid transverse magneto-thermoelectric conversion has been experimentally demonstrated in several systems. In Ref. [98], ATML composed of Bi88Sb12, which exhibits the large magneto-thermoelectric effects, and BST, which exhibits the large Seebeck and Peltier effects, was used to enhance the transverse thermoelectric conversion performance by applying a magnetic field. The observed performance enhancement in this system originates from the hybrid thermoelectric conversion involving three effects: the off-diagonal Seebeck (Peltier), magneto-Seebeck (Peltier), and ordinary Nernst (Ettingshausen) effects [Fig. 10(a)]. Figure 10(b) shows the charge current $J$ dependence of the surface temperature averaged over the Bi88Sb12/BST junction in ATML under the steady state, where the contribution of Joule heating results in a parabolic behavior and that of the thermoelectric effects results in the shift of the parabolic curves. The systematic experimental results reported in Ref. [98] indicate that a temperature gradient is generated in a direction perpendicular to the applied charge current and the steady-state cooling performance due to the transverse thermoelectric conversion is dramatically modulated by the reversal of a magnetic field. In the experimental setup depicted in Fig. 10(b), when a positive magnetic field is applied, the direction of the net heat current generated by the off-diagonal Peltier effect is the same as that by the ordinary Ettingshausen effect, improving the transverse thermoelectric conversion performance. The figure of merit at room temperature for Bi88Sb12/BST ATML is 0.20 (0.14) when a positive (negative) magnetic field is applied, and this field-dependent modulation is much larger than the figure of merit for ONE alone in the Bi88Sb12 alloy. This is attributed to the contribution of the cross term in Eq. (102). Furthermore, in this study, it has been demonstrated that by replacing BST in ATML with the permanent magnet Nd2Fe14B, the magneto-thermoelectric effects can be superimposed without applying an external magnetic field due to the stray magnetic field from the remanent magnetization of Nd2Fe14B. In Ref. [99], the candidate phenomena for the hybrid transverse magneto-thermoelectric conversion have been expanded by inducing ANE in ATML comprising the topological ferromagnet Co2MnGa. In Co2MnGa-based ATMLs combined with $p$-type BST and with $n$-type Bi2Te3, a magnetization-direction-dependent modulation of the figure of merit significantly larger than that of ANE alone was observed owing to the hybridization of the off-diagonal Seebeck effect and ANE.
@@
Examples of the hybrid transverse magneto-thermoelectric conversion are still limited. In principle, the Seebeck-driven ordinary/anomalous Hall effects and spin-Seebeck/spin-dependent-Seebeck-driven ISHEs, not used in the above examples, are also integrable into the hybrid transverse magneto-thermoelectric conversion. In fact, although the thermoelectric output is small, the hybrid transverse thermoelectric generation based on the combination of ANE and spin-Seebeck-driven ISHE has been demonstrated [83, 85, 147]. Originally, spin-caloritronic phenomena such as ANE and SSE have been extensively studied from the viewpoint of fundamental physics, but their low thermoelectric conversion performance has limited their impact on applications. The concept of the hybrid transverse magneto-thermoelectric conversion provides one of the solutions for practical applications of such spin-caloritronic phenomena.

### v2

@@ V.2 Effect of thermal boundary conditions
Figure 7: (a),(b) Schematics of the measurement setups for the transverse thermoelectric generation in the nearly isothermal and adiabatic conditions. (c) Schematic of the measurement setup for the transverse thermopower in SmCo5-type permanent magnet slabs with different values of the width $d$. (d) $H$ dependence of the transverse electric field $V_{y}/d$ for the SmCo5 slabs with $d=7.7~{\rm mm}$ (red and orange) and 2.1 mm (black and gray) at $|\nabla_{x}T|\sim 1\times 10^{3}~{\rm K/m}$. The $H$-independent offset of the transverse electric field is subtracted from the raw data. The red and black (orange and gray) curves show the $H$ dependence of $V_{y}/d$ after (before) subtracting the $H$-linear component estimated by linear fitting of the signals in the high $H$ region. (e) $|\nabla_{x}T|$ dependence of the $M$-odd-dependent component of the transverse electric field $V_{\rm odd}/d$, which is estimated by extrapolating the $V_{y}/d$ signal in the high $H$ region to zero field.
@@
We now experimentally demonstrate the influence of sample dimensions on the transverse thermopower using a bulk polycrystalline SmCo5 slab. The SmCo5 block, commercially available from Magfine Corporation, Japan, was cut into two rectangular slabs with dimensions along the $x$, $y$, and $z$ directions respectively being 10.0, 7.7, and 0.8 mm and 10.0, 2.1, and 0.8 mm using a diamond wire saw, where only the length along the $y$ direction, $d$, is significantly different. The slabs have an easy axis of magnetization along $z$ direction. The measurement configuration is illustrated in Fig. 7(c). The two SmCo5 slabs were bridged between two heat baths separated by 8.0 mm and firmly attached to the heat baths. To apply a temperature difference between the heat baths, chip heaters were attached to the hot side. The sample holder was fixed at the center of a superconducting magnet, where a magnetic field ranging from $\mu_{0}H=5~{\rm T}$ to $-5~{\rm T}$ was applied along the $z$ direction. The $H$ dependence of the transverse voltage $V_{y}$ along the $y$ direction was recorded by attaching Cu wires at the ends of the slabs at various values of the temperature difference. To accurately measure the temperature gradient on the SmCo5 slabs, the top surface of the slabs was coated with insulating black ink with an emissivity of >0.94>0.94> 0.94 and steady-state temperature images were captured using an infrared camera.
@@
Figure 7(d) shows the $H$ dependence of the transverse electric field $V_{y}/d$ for the SmCo5 slabs, recorded under a longitudinal temperature gradient of $|\nabla_{x}T|\sim 1\times 10^{3}~{\rm K/m}$. In both the slabs, the transverse electric field changed its sign with large hysteresis with respect to the $H$ reversal. The $H$ dependence of $V_{y}/d$ corresponds to the magnetization curve of the SmCo5 permanent magnet, indicating the appearance of ANE. To determine the magnitude of the $M$-dependent transverse thermopower for the SmCo5 slabs, the $M$-odd-dependent component of $V_{y}/d$ under the saturation state, $V_{\rm odd}/d$, was extracted for each $|\nabla_{x}T|$ value by extrapolating the signal in the high $H$ region to zero field. Here, the finite $H$-linear component was observed only for the slab with $d=2.1~{\rm mm}$, suggesting the presence of the Seebeck voltage between the SmCo5/Cu-wire contacts induced by the ordinary ($H$-dependent) THE in SmCo5 [Fig. 7(d)]. Although the SmCo5 slabs were cut from the same ingot, they exhibited different $V_{\rm odd}/d$ values [Fig. 7(e)]. The $M$-dependent transverse thermopower for the SmCo5 slab with $d=2.1~{\rm mm}$ ($d=7.7~{\rm mm}$) was found to be $5.1~\mu{\rm V/K}$ ($4.1~\mu{\rm V/K}$); the difference in the transverse thermopower is much larger than measurement errors. Recall the fact that the slab with $d=2.1~{\rm mm}$ ($d=7.7~{\rm mm}$) is closer to the adiabatic (isothermal) condition. Therefore, the transverse thermopower under the adiabatic condition is expected to be overestimated compared to that under the isothermal condition because of the same sign of $S_{xy}$ and the correction term induced by the Seebeck effect and anomalous ($M$-dependent) THE. This expectation is consistent with the observed difference in the $M$-dependent transverse thermopower. Here, we note that the correction term discussed theoretically in Sec. III is determined by the absolute Seebeck coefficient of a magnetic material (i.e. SmCo5), while the observed difference in the transverse thermopower in Fig. 7 depends on the relative Seebeck coefficient between SmCo5 and Cu. However, since the Seebeck coefficient of Cu ($<2~{\rm\mu V/K}$) is much smaller than that of SmCo5 ($-19~{\rm\mu V/K}$ [63]) at room temperature, the magnitude and sign of the correction term in the present setup is determined mainly by the transport properties of SmCo5. The experimental results shown here offer guidelines for determining the appropriate sample aspect ratio and measurement setup for transverse thermopower measurements, minimizing parasitic signals from the Seebeck effect and THE. We also note that considering typical sample dimensions, the estimation of $S_{\rm ANE}$ from the measurement of the anomalous Ettingshausen effect and the Onsager reciprocal relation reduces the influence of the correction term [29].
@@ VI.1 Heat flux sensing
A heat flux sensor is a device enabling simultaneous detection of the magnitude and direction of a heat flow, which can be an essential component for smart thermal management systems. However, practical applications of heat flux sensors have been limited due to several problems. The commercially available sensors are based on the Seebeck effect and consist of a serially connected three-dimensional array of two different thermoelectric materials. Therefore, the heat flux sensors based on the Seebeck effect require a durable substrate or thick plate to provide mechanical stability. Hence, the conventional sensors are mainly applicable to flat surfaces and their mechanical flexibility is limited. Due to the presence of thick substrates, these sensors have large thermal resistance that changes the heat flow distribution to be detected. Although the sensibility of the heat flux sensors based on the Seebeck effect is proportional to the number of the thermoelectric material junctions and to the sensor size, the complex structure and low mechanical durability make it difficult to construct highly sensitive sensors. As an alternative to the sensors having these drawbacks, thin-film-based heat flux sensors driven by the transverse thermoelectric effects, e.g. SSE-driven ISHE [160] and ANE [37, 48, 161, 162], were proposed and demonstrated. Owing to its simpler structure and improved transverse thermopower, ANE-based heat flux sensors have recently become mainstream. In the ANE-based sensor, a lateral thermopile structure consisting of alternately arranged and serially connected two different wires with different $S_{\rm ANE}$ values is used. The simple thermopile structure and the symmetry of ANE make the sensors thin, reducing their thermal resistance. If magnetic materials showing large ANE can be formed on thin flexible sheets, flexible heat flux sensors can be constructed [48, 161]. These features of the ANE-based heat flux sensors are advantageous over the conventional heat flux sensors based on the Seebeck effect.
Figure 8: (a) Schematic of the ANE-based heat flux sensor. The sensor comprises the amorphous Sm20Co80 wires and Au wires arranged alternately and connected in series formed on a flexible polyethylene naphthalate (PEN) sheet. ${\bf q}$ denotes the heat flux density with the magnitude $q$. (b) $H$ dependence of $V$ for the Sm20Co80/Au sensor for various values of $q$. (c) $q$ dependence of $V_{\rm odd}$ at $|\mu_{0}H|=0.2$ and 0.0 T. The data in (b) and (c) are reconstructed from Fig. 6 of Ref. [161].
@@
[…]
As shown in Fig. 8(b), clear ANE signals with hysteresis behaviors were observed in the $H$ dependence of $V$ and their magnitude increases with increasing $q$ across the device. The saturated $V$ values for $\mu_{0}H>0.1~{\rm T}$ indicate the negligibly small ONE contribution, enabling the simple extraction of the ANE contribution. Figure 8(c) shows the ANE voltage, represented by $V_{\rm odd}$, as a function of $q$ [161], where $V_{\rm odd}$ at 0.2 T was extracted by the field-odd component of $V$ at $|\mu_{0}H|=0.2~{\rm T}$ and $V_{\rm odd}$ at zero field was extracted from the zero-field $V$ values for the positive-to-negative and negative-to-positive sweep of $H$. The sensitivity of the ANE-based heat flux sensor was estimated to be $V_{\rm ANE}/q\sim V_{\rm odd}/j_{\rm q}=1.12\pm 0.02\times 10^{-7}~{\rm VW}^{-1}{\rm m}^{2}$ and $0.89\pm 0.02\times 10^{-7}~{\rm VW}^{-1}{\rm m}^{2}$ in the saturation magnetization (0.2 T) and remanent magnetization (0.0 T) states, respectively. Importantly, the use of magnetic thin films with finite coercivity and remanent magnetization makes it possible to sense a heat flux using ANE even in the absence of an external magnetic field. Therefore, the design and control of magnetic anisotropy are essential for the application of ANE. Although the sensitivity of the sensor used in this demonstration is lower than that of commercial sensors based on the Seebeck effect, ANE has strong advantages in flexibility and low thermal resistance, extending applications of heat flux sensors. To further improve the performance of the ANE-based heat flux sensors, extensive efforts have been devoted to developing new materials with larger $S_{\rm ANE}$ and lower $\kappa$ as well as to optimizing the device structure. Since amorphous materials can be fabricated on any surface and have lower $\kappa$ than crystalline materials, they are promising candidates for the development of low-cost, high-performance heat flux sensors with large areas [59, 60, 61, 62, 161].
@@ VI.2 Thermal energy harvesting
Transverse thermoelectric devices are useful for harvesting and utilizing widely distributed thermal energy. By installing simple block- or sheet-shaped transverse thermoelectric devices on a heat source surface, one can extract a charge current and voltage along the surface direction from a temperature gradient perpendicular to the surface. Kirihara et al. pointed out that SSE-driven ISHE is suitable for the thermal energy harvesting from curved or complex-shaped heat source surfaces and demonstrated that SSE devices can be fabricated using simple coating [90] or spray [160] processes. Similar functionality can be obtained using ONE (ANE) in the in-plane magnetic field (magnetized) configuration, enabling the fabrication of flexible transverse thermoelectric devices based on thin films deposited on flexible substrates [161] or amorphous metal ribbons [59, 60, 61, 62]. However, challenges include the difficulty of extracting a large charge current and electrical power from thin films and the fact that the thermopower generated by ONE, ANE, and SSE-driven ISHE has not yet reached sufficient levels for practical applications.
@@
Experimental demonstrations of thermal energy harvesting using the off-diagonal Seebeck effect, which exhibits high transverse thermoelectric conversion performance, have also been reported. Figure 9 compares the maximum power density normalized by the square of the temperature gradient ($\propto$ the power factor) for thermoelectric conversion modules driven by various principles [4, 9, 10, 169, 170, 171, 172, 99, 100, 101, 167, 108]. The maximum power density of the ATML-based modules driven by the off-diagonal Seebeck effect stands out, being 2-3 orders of magnitude larger than that of the modules driven by ONE and ANE. Recently developed SmCo5/BST ATML exhibits excellent transverse thermoelectric conversion performance, with the maximum power density of the modules composed of 14 ATML elements reaching 56.7 mW/cm2 at a temperature difference of 152 K [108]. SmCo5/BST ATML has higher thermal conductivity than longitudinal thermoelectric materials, making it difficult to maintain a large temperature difference in steady-state conditions; the efficiency of the ATML-based module falls short of that of longitudinal modules driven by the Seebeck effect. However, the maximum power density normalized by the square of the temperature gradient in ATMLs exceeds that of the longitudinal modules based on the Seebeck effect (Fig. 9). Therefore, ATML-based transverse thermoelectric modules are suitable for thermal energy harvesting because they can generate large power even in environments where only small temperature differences can be generated. Additionally, in transverse thermoelectric modules with thermopile structures, by alternately changing the height of each element, a built-in heat sink can be incorporated to promote heat release to the atmosphere [108]. Such a built-in heat sink structure cannot be achieved in longitudinal modules where rigid plates are present on the top and bottom surfaces. Even without sandwiching the thermoelectric module between hot and cold heat baths, the built-in heat sink structure enables substantially large power generation simply by placing it on a heat source owing to the improved heat release to the atmosphere. In particular, SmCo5/BST ATML has remanent magnetization and functions as a permanent magnet; when it can be attached to a heat source through a magnetic attractive force, the contact thermal resistance between the module and heat source decreases even when thermal interface materials are inserted between them, enabling the effective generation of a temperature difference from the heat source. As shown in Ref. [108], when the thermopile module comprising SmCo5/BST ATML elements is placed on a hot plate, the electrical power generated solely by heat release to the air reaches a value sufficient to drive small sensors.
@@ VI.3 Hybrid transverse magneto-thermoelectric conversion
In this subsection, we introduce one of the promising strategies for improving transverse thermoelectric conversion performance: hybrid transverse magneto-thermoelectric conversion [106, 107, 109]. As classified in Fig. 3, there are various driving principles for the transverse thermoelectric conversion, but the phenomena with time-reversal symmetry breaking and those with structural symmetry breaking have been studied independently in different fields. The interdisciplinary fusion of these phenomena not only leads to the creation of new research fields but also significantly contributes to improving the transverse thermoelectric conversion performance. Here, imagine a situation where multiple transverse thermoelectric conversion principles manifest in a single material, i.e. either a homogeneous or composite/hybrid material. As an example, if the transverse thermopower due to the off-diagonal Seebeck effect $S_{\rm OD}$ and that due to ONE $S_{\rm ONE}$ are simultaneously manifested, the total transverse thermopower is simply the sum of these two contribution: $S_{\rm OD}+S_{\rm ONE}$. Since $|S_{\rm OD}|>|S_{\rm ONE}|$ in many cases, the contribution of ONE to the total transverse thermopower is small. In this situation, the figure of merit is proportional to the square of the total transverse thermopower, and described as
@@
The effectiveness of the hybrid transverse magneto-thermoelectric conversion has been experimentally demonstrated in several systems. In Ref. [106], ATML composed of Bi88Sb12, which exhibits the large magneto-thermoelectric effects, and BST, which exhibits the large Seebeck and Peltier effects, was used to enhance the transverse thermoelectric conversion performance by applying a magnetic field. The observed performance enhancement in this system originates from the hybrid thermoelectric conversion involving three effects: the off-diagonal Seebeck (Peltier), magneto-Seebeck (Peltier), and ordinary Nernst (Ettingshausen) effects [Fig. 10(a)]. Figure 10(b) shows the charge current $J$ dependence of the surface temperature averaged over the Bi88Sb12/BST junction in ATML under the steady state, where the contribution of Joule heating results in a parabolic behavior and that of the thermoelectric effects results in the shift of the parabolic curves. The systematic experimental results reported in Ref. [106] indicate that a temperature gradient is generated in a direction perpendicular to the applied charge current and the steady-state cooling performance due to the transverse thermoelectric conversion is dramatically modulated by the reversal of a magnetic field. In the experimental setup depicted in Fig. 10(b), when a positive magnetic field is applied, the direction of the net heat current generated by the off-diagonal Peltier effect is the same as that by the ordinary Ettingshausen effect, improving the transverse thermoelectric conversion performance. The figure of merit at room temperature for Bi88Sb12/BST ATML is 0.20 (0.14) when a positive (negative) magnetic field is applied, and this field-dependent modulation is much larger than the figure of merit for ONE alone in the Bi88Sb12 alloy. This is attributed to the contribution of the cross term in Eq. (102). Furthermore, in this study, it has been demonstrated that by replacing BST in ATML with the permanent magnet Nd2Fe14B, the magneto-thermoelectric effects can be superimposed without applying an external magnetic field due to the stray magnetic field from the remanent magnetization of Nd2Fe14B. In Ref. [107], the candidate phenomena for the hybrid transverse magneto-thermoelectric conversion have been expanded by inducing ANE in ATML comprising the topological ferromagnet Co2MnGa. In Co2MnGa-based ATMLs combined with $p$-type BST and with $n$-type Bi2Te3, a magnetization-direction-dependent modulation of the figure of merit significantly larger than that of ANE alone was observed owing to the hybridization of the off-diagonal Seebeck effect and ANE. In Ref. [109], the hybridization of the off-diagonal Seebeck effect and ANE has been realized in SmCo5/BST ATML in the absence of an external magnetic field by utilizing the large in-plane remanent magnetization of SmCo5. Note that in these ATMLs, the transverse thermopower generated by ONE (ANE), $S_{\rm ONE}$ ($S_{\rm ANE}$), is effectively reduced because of the shunting effect in adjacent conductors. Nevertheless, the gain from the cross term in Eq. (102) is greater than the reduction in $S_{\rm ONE}$ ($S_{\rm ANE}$).
@@
Examples of the hybrid transverse magneto-thermoelectric conversion are still limited. In principle, the Seebeck-effect-driven ordinary/anomalous Hall effects and SSE/SdSE-driven ISHEs, not used in the above examples, are also integrable into the hybrid transverse magneto-thermoelectric conversion. In fact, although the thermoelectric output is small, the hybrid transverse thermoelectric generation based on the combination of ANE and SSE-driven ISHE has been demonstrated [91, 93, 158]. Originally, spin-caloritronic phenomena such as ANE and SSE have been extensively studied from the viewpoint of fundamental physics, but their low thermoelectric conversion performance has limited their impact on applications. The concept of the hybrid transverse magneto-thermoelectric conversion provides one of the solutions for practical applications of such spin-caloritronic phenomena.